Draw the Lewis structure for SeS3 and answer the following questions.

How many valence electrons are present in this compound?

How many bonding electrons are present in this compound?

How many lone pair (non-bonding) electrons are present in this compound?
Answer #1
Concepts and reason

Lewis structure gives the detailed explanation on connectivity of atoms present in the molecule and also indicates the lone pair electron of atoms in the molecule.

Valence electrons: They are the electrons in the outermost shell of an atom, which can participate in bonding.

Octet rule: According to octet rule, when an atom contains eight electrons in its valence shell (outermost shell), it will be stable. Most of the atom will donate or accept electrons to its outer most orbital to satisfy the octet rule.

Fundamentals

The steps to draw Lewis structure are as follows:

Step 1: Find the number of valence electrons.

Step 2: Find the number of electrons needed to make the atoms complete its octet.

Step 3: Determine the number of bonds in the molecule.

Step 4: Choose a central atom.

Step 5: Draw a skeletal structure.

Step 6: Place electrons around the atoms.

The example for Lewis structure:

H

The water molecule contains one oxygen atom and 2 hydrogen atoms. The valance electron of oxygen atom is 6 and hydrogen atom is 1.

Therefore, in total, there are 6+2(1)=8
electrons. In this molecule, oxygen is bonded to two hydrogen atoms. Therefore, for 2 bonds, 4 electrons are contributed in the formation of these bonds and the remaining 4 electrons remains as non-bonded electrons which are indicated as lone pair electrons on oxygen atom.

The given compound is,.

The Lewis structure of is as follows:

=

The given compound is,.

The number of valence electrons in sulfur is

The number of valence electrons in selenium is

Total number of valence electrons = Se +(3xS)
= 6+(36)
= 24

Using the Lewis structure of ,find the bonding electrons.

S:
Number of bonds = 4

Total number of bonding electrons = 2 x Number of bonds
= 2 x 4
= 8

Therefore, the total number of bonding electrons is.

By using the Lewis structure of, find the lone pair electrons.

:S-
:00:
Se-
S:
5:
Number of lone pairs = 8

Total number of lone pair electrons = 2 x Number of lone pairs
= 2x8
= 16

Therefore, the total number of lone pair (non-bonding) electrons is.

Ans:

The total number of valence electrons present in this compound is.

The total number of bonding electrons present in this compound is.

The total number of lone pair (non-bonding) electrons present in this compound is.